Operations Report – May 16th

Crew 212 Operations Report 16-MAY-2019

SOL: 11

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None.

Notes on non-nominal systems: Nothing to report.

Generator (hours run last night): [From Director]
On (last night): 6:15 PM
Off (this morning): [From Director]

Solar— SOC 84% at 7:40 PM

Diesel Reading – 82%

Propane Reading – 45%

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 200 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 113 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143754.6

Toilet tank emptied: No

Deimos rover used: no (out of order)
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes [USED BY DIRECTOR]
Hours: [From Director]
Beginning charge: 100%
Ending charge: 71%
Currently charging: Yes

Opportunity rover used: Yes [USED BY DIRECTOR]
Hours: [From Director]
Beginning charge: 76%
Ending charge: 60%
Currently charging: Yes

Curiosity rover used: No
Hours: 96.6h
Beginning charge:
Ending charge:
Currently charging: Yes

Notes on rovers: [From Director] Sojo’s battery bank was cleaned and batteries refilled with distilled water. Only one cell was critically low (below the metal) but it was still wet. 1.5 gallons of distilled water used. No connections were loose and no battery caps were trapped under the security bar. The thermistor on the Rangers need to go to get serviced. All the batteries on Oppi were cleaned and the batteries will be filled tomorrow. Oppi needs the brakes serviced, as they have failed and run out of brake fluid twice this season.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: after EVA#11, we had a debriefing and we worked on our experiments. The high winds have slightly damaged the tunnel covers in a few points.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: the new pump is working fine, but I just wanted a clarification: is it normal that the pump works in a pulsed mode (oscillating ON/OFF) when a low quantity of water is required?

EVA #12 Report – May 16th

Crew 212 EVA Report 17-MAY-2019

EVA #012

Author: Zoe Townsend

Purpose of EVA: Collection of sand for outreach purposes and physiological experiments to see how the suits performed in exercise.

Start time: 09.30

End time: 11.00


Today’s EVA was aimed at assisting with the project on the physiological effects of the suits. The EVA was to walk up to the Mid Ridge Planitia viewpoint to see how much oxygen was in the blood before and after the EVA, therefore assessing the suitability of the suits.

The experiment was deemed a partial failure where there were difficulties in isolating the suit to the outside elements. The windy weather also affected the testing where it wasn’t a true test of the suits fans ability.

Summary of the debrief, after the EVA:

  • · Communication was the best over all the sim
  • · Issue with clarification of time change (read earlier email from commander)

Coordinates (UTM NAD27 CONUS): Sector12S Northern 4251250, Eastern 516000

EVA Participants: Paola, Mariona (Medic), Camilo and Zoe (Commander)

Road(s) and routes per MDRS Map: Sagan Street 1103

Mode of travel: (walking or driving)? Walking.

Greenhab Report – May 16th

Crew 212 GreenHab Report – 16-May-2019: SOL 11

GreenHab Officer: Hector Palomeque

Environmental control: Ambient with the door open from 07:00 to 12:00 + fan turned up to 3 (max) from 11:00 to 19:00

Shade cloth 80%: on

Average temperatures:

Low: 21 °C

High: 41 °C

Hours of supplemental light: 0

Daily water usage for crops: 9 gallons

Water in Blue Tank: 113 gallons

Time(s) of watering for crops: 07:30

Changes to crops: All plants look healthy. Basil, tomato and cilantro plants continue blooming. The yellow tomatoes remaining now are ready to harvest. Chard is completely recovered from the previous harvest. Sugar Snap Peas are growing better under shade (more than 80%). Tomato in the big blue tray continue with its fructification, now it has 3 tomatoes.

Narrative: The GreenHab fan had to be turned on because of the indoor temperature has reached +40°C at 11:00. The GreenHab door had to be closed at 12:00 due to the strong winds. All plants were watered once early in the morning to take advantage of the fresh environment.

Harvested: Tomato 97g + Basil 36g

Support/supplies needed: None.

Spacesuit Report – May 16th

Crew 212-LATAM-III

Submitted by Spacesuit Engineer: Camilo Andrés Zorro Mendoza

Spacesuit report

Sol: 11

Sim suit used Engineering EVA: 7, 8, 5 and 4.
used for 1 hour and half

Sim suit: 7
Voltage level before EVA: 12.8 V
Voltage level after EVA:
12.5 V

Sim suit: 4
Voltage level before EVA:
13.1 V
Voltage level after EVA:
12.6 V

Sim suit: 5
Voltage level before EVA: 12.6 V
Voltage level after EVA:
5.6 V (it will be changed)

Sim suit: 8
Voltage level before EVA:
13 V
Voltage level after EVA:
12.8 V

Journalist Report – May 16th

Today was Sol 11.

We undertook a morning EVA, the idea was the test the suits for the suitability whilst undertaking exploration of another planet. This was aimed at looking at oxygen in the blood after exercise, therefore the EVA was walking up to a ridge such as would be typical for explorations in the future.

Afterwards the crew members continued on with their individual’s projects. There was progress in processing the Eastern Veil Nebula photos, soy bacteria was grown in petri dishes this is for a logically first step identification method, in the radiation bacteria project 1 lettuce has sprouted and finally, the rover revival has continued where now 2 out of 4 wheels are functional, there is hope that it will make one EVA before the end of sim.

Due to poor weather most of the Crew stayed within the Hab today. Therefore, the group made cookies and prepped the pizza dough for Shannon’s arrival tomorrow.

Sol Summary – May 16th

Crew 212 Sol 11 Summary Report 16-MAY-2019

Sol: 11

Summary Title: Martian Lessons

Author’s name: Marlen Castillo Vilcahuaman

Mission Status: A-ok

Sol Activity Summary: Today was a windy day for EVAs. After asking permission to the Director (with the previous anouncement to Mission Support after an issue with pasta beating), some crew members went to the walking EVA. We’re happy to notice that, today, we kept communication all the time. Radios where performing fine and the Hab could always communicate with the EVA crew.

Look Ahead Plan: Trying to use the most of our last days in MDRS. EVA requests for tomorrow.

Anomalies in work: None.

Weather: Windy, really windy.

Crew Physical Status: A-ok

EVA: 1

Reports to be file: EVA report, Space Suit Report, Operations report, Green Hab report, Journalist report, Daily Pictures, EVA requests

Support Requested: There’s a delivery for Saturday for one of the crew members. They have sent it to the PO Box.

Copyright © The Mars Society. All rights reserved. | Main Site